49. Airline passengers arrive randomly and independently at the passenger-screening facility at a major international airport. The mean arrival rate is 10 passengers per minute.

PLEASE SHOW ANSWERS AND FORMULAS IN EXCEL

a. Compute the probability of no arrivals in a one-minute period.

b. Compute the probability that three or fewer passengers arrive in a one-minute period.

c. Compute the probability of no arrivals in a 15-second period.

d. Compute the probability of at least one arrival in a 15-second period.
Answer #1


Part a)

Excel formula POISSON(0,10,FALSE)

Part b)




P ( X <= 3 ) = 0.0103

Excel formula POISSON(3,10,TRUE)


Part c)

Excel formula  POISSON(0,4,FALSE )

Part d)

P ( X >= 1 ) = 1 - P ( X = 0 ) = 1 - 0.0183 = 0.9817

Excel formula 1 -  POISSON(0,4,FALSE )